Exercise 11-9 Mixed cost at different levels of activity LO 11-1 Stuart Corporation paid one of its sales representatives $7,000 during the month of March. The rep is paid a base salary plus $11 per unit of product sold. During March, the rep sold 150 units. Required Calculate the total monthly cost of the sales representatives salary for each of the following months: Month AprilMayJuneJuly 190 200 110 Number of units sold Total variable cost Total fixed cost Total salary cost 100 O S
Answer #1
Total salary for march $ 7,000
Less: Variable cost ($11*150) $ 1,650
Fixed cost per month $ 5,350
Month April May June July
Number of unit sold 190 100 200 110
Total variable cost $11*190 = $2,090 $11*100 = $1,100 $11*200 = $2,200 $11*110 = $1,210
Total fixed cost $                   5,350 $                   5,350 $                   5,350 $                   5,350
Total salary cost $                   7,440 $                   6,450 $                   7,550 $                   6,560

You can reach me over comment box if you have any doubts. Please rate this answer